# PAC-Chernoff Bounds: Understanding Generalization in the Interpolation Regime

### v3

@@ 4 The Rate Function Characterizes the Generalization of Interpolators
In this section, we showcase the main results of this work, Theorems 4.1 and 4.5, which show how the (inverse) rate function can characterize the generalization performance of interpolators. Furthermore, we present a new notion of smoothness of a model which perfectly aligns with the following statement: a (sufficiently) smoother interpolator generalizes better.
@@ 4.1 Tight Distribution-Dependent Bounds for Over-parameterized Interpolators
As discussed in the introduction, current high-probability generalization bounds in the form of Equation (1) have been unable so far to explain the generalization of over-parameterized interpolators. Different works (Nagarajan and Kolter, 2019c; Nagarajan, 2021; Gastpar et al., 2023; Wang et al., 2024) have also shown that the problem lies in the impossibility of having tight generalization bounds solely depending on the training data for over-parameterized model classes. Gastpar et al. (2023) even concludes that “a bound without explicit distributional assumptions is likely to be not tight”. Thus, an open question raises:
@@
The answer to this question does not seem to be straightforward given that Nagarajan and Kolter (2019c) also showed examples of distribution-dependent bounds (Nagarajan and Kolter (2019c)’s Definition 3.3) which are not tight. However, Chernoff bounds provide an answer to the above question. The following result shows an uniform-convergence bound which applies simultaneously over the model class,
@@
Going back to Theorem 4.1, in the above bound, ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\textstyle\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)$ defines a complexity measure for the model ${\boldsymbol{\theta}}$ in the context of a model class defined by $p$ parameters, the training data set of $n$ samples and, according to the own definition of ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\cdot\right)$ (see Definition 2), it also depends on the data-generating distribution $\nu$. From Proposition 3.2, this complexity measure ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\textstyle\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)$ monotonically grows with the size of the model class and monotonically decreases with the level of confidence $\delta$ and the size of the training data. Actually, the following result shows that, when $n$ goes to infinity, the proposed complexity measure converges with rate $1/\sqrt{n}$ to the (scaled) standard deviation of the loss function.
@@
However, the relevant property of this complexity measure is that it is a perfectly tight proxy of the expected loss $L({\boldsymbol{\theta}})$ for models that interpolate the training data, $\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon$, even if the model class is over-parameterized,
@@
In this case, the event that holds with probability $1-\delta$ is the implication in Equation 3.
Using there results, we can derive an algorithm-distribution-dependent bound that is perfectly tight for any data distribution and any learning algorithm; even for over-parameterized model classes. However, for this bound to be tight it is imperative that the model interpolates the training data. Formally, for any fixed $\epsilon>0$, with $2^{\mathcal{X}\times\mathcal{Y}}$ denoting the power set of all possible data sets of any size; let $\mathcal{A}_{\epsilon}:2^{\mathcal{X}\times\mathcal{Y}}\to{\boldsymbol{\Theta}}$ be an algorithm that takes any data set and returns a model ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}$ such that its training loss is lower than $\epsilon$. That is, for any training data set $D\in 2^{\mathcal{X}\times\mathcal{Y}}$, $\mathcal{A}_{\epsilon}$ verifies that ${\hat{L}}({\cal A}_{\epsilon}(D),D)\leq\epsilon$. This kind of algorithms are quite common in machine learning where the model space ${\boldsymbol{\Theta}}$ is big enough that there exists models capable of memorizing random labels (Zhang et al., 2017, Theorem 1). Then, for any ${\cal A}_{\epsilon}$ algorithm, with h.p. $1-\delta$ over $D\sim\nu^{n}$,
@@
The above result provides an answer to the Open Question 4.1:
@@ 4.2 Smoother interpolators generalize better
This result showcases the relation between the smoothness and the variance the loss function between a pair of models ${\boldsymbol{\theta}}$ and ${\boldsymbol{\theta}}^{\prime}$. As a result, we can intuitively understand this notion of smoothness as a generalization of the variance of the loss of a model. Using this definition of smoothness and the results presented in Section 4.1, we can easily derive a result providing an answer to which interpolator generalizes better,
@@
If we look again at Figure 2 and assuming that the definition of smoothness given in 3 applies to the estimated rate functions, one can easily see that Theorem 4.5 perfectly explains when a interpolator generalizes better than other.
From the results of the current and the previous section, it is clear that the rate function of an interpolator is the key quantity defining its generalization error. The question now is whether this theoretical analysis based on the use of distribution-dependent quantities is useful or not to understand the inner workings of current learning techniques.
@@ 5 Explicit Regularization
Structural risk minimizing (Shawe-Taylor et al., 1998) is a learning principle based on regularizers that penalize the complexity of a model when minimizing the training loss. If we denote $r({\boldsymbol{\theta}})$ to a regularizing function, the learning objective should be the following,
@@
High-probability bounds of the form given in Equation (1), $L({\boldsymbol{\theta}})\leq\hat{L}(D,{\boldsymbol{\theta}})+{\cal C}({\boldsymbol{\theta}},D,\delta)$, have been traditionally used to justify the use of regularizers and to derive novel ones. According to these bounds, an obvious choice for regularizes is the complexity term, $r({\boldsymbol{\theta}})={\cal C}({\boldsymbol{\theta}},D,\delta)$. Then, the structural risk minimization problem given in Equation (4) would correspond to minimize a high-probability upper bound over the expected loss $L({\boldsymbol{\theta}})$. Many regularizing methods have been studied from this perspective. One of the most commons one is being the $\ell^{2}$-norm regularization and other variations based on spectral norms, which recurrently show up within the complexity term of many generalization bounds (Bartlett et al., 2017). However, as discussed in the introduction, these bounds are known to be loose in the context of over-parameterized model classes interpolating the data. In consequence, the rationale of using their associated complexity measures ${\cal C}({\boldsymbol{\theta}},D,\delta)$ as regularizes is no longer as strong as it used to be in learning setups where the model class is not over-parameterized.
Open Question 3 Is there a regularizer $r({\boldsymbol{\theta}})$ that ensures near-optimal performance for over-parameterized model classes interpolating the training data?
The following result provides an answer to the Open Question 5. It shows that the complexity measure of the complexity measure of the PAC-Chernoff bound given in Theorem 4.1 can be used as a regularizer. That is to say, using the inverse rate function ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)$, we get an optimal regularizer for over-parameterized interpolators. More precisely, for any $\epsilon>0$, let ${\boldsymbol{\theta}}^{\star}_{\epsilon}$ and ${\boldsymbol{\theta}}^{\crossproduct}_{\epsilon}$ denote the interpolator with the best generalization performance and the interpolator with the smallest inverse rate function, respectively,
@@
More intuitively, promoting models with a small inverse rate function is equivalent to promoting models with a larger rate function or, equivalently, smoother models. In that sense, the above result shows that the smoothest interpolator is also a near-optimal interpolator in terms of generalization performance.

### v4

@@ 4 The Rate Function Characterizes the Generalization of Interpolators
In this section, we showcase the main results of this work, Theorems 4.1 and 4.6, which show how the (inverse) rate function can characterize the generalization performance of interpolators. Furthermore, we present a new notion of smoothness of a model which perfectly aligns with the following statement: a (sufficiently) smoother interpolator generalizes better.
@@ 4.1 Tight Distribution-Dependent Bounds for Over-parameterized Interpolators
As discussed in the introduction, current high-probability generalization bounds in the form of Equation (1) have been unable so far to explain the generalization of over-parameterized interpolators. Different works (Nagarajan & Kolter, 2019b, ; Nagarajan,, 2021; Gastpar et al.,, 2024; Wang et al.,, 2024) have also shown that the problem lies in the impossibility of having tight generalization bounds solely depending on the training data for over-parameterized model classes. Gastpar et al., (2024) even concludes that “a bound without explicit distributional assumptions is likely to be not tight”. Thus, an open question raises:
@@
The answer to this question does not seem to be straightforward given that Nagarajan & Kolter, 2019b also showed examples of distribution-dependent bounds (Nagarajan & Kolter, 2019b ’s Definition 3.3) which are not tight. However, the following result shows an uniform-convergence bound which applies simultaneously over the model class,
@@
Going back to Theorem 4.1, in the above bound, ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\textstyle\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)$ defines a complexity measure for the model ${\boldsymbol{\theta}}$ in the context of a model class defined by $p$ parameters, the training data set of $n$ samples and, according to the own definition of ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\cdot\right)$ (see Definition 2), it also depends on the data-generating distribution $\nu$. From Proposition 3.2, this complexity measure ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\textstyle\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)$ monotonically grows with the size of the model class and monotonically decreases with the level of confidence $\delta$ and the size of the training data.
It is common to find models, within the same model class, that define the same loss function, $\ell(\cdot,\cdot,{\boldsymbol{\theta}}_{1})=\ell(\cdot,\cdot,{\boldsymbol{\theta}}_{2})$. For instance, in a Multi-Layer Perceptron (MLP), weights can be permuted in specific ways without altering model predictions, and models with zeros in the last layer produce the same predictions regardless of the weights in other layers. Consequently, when calculating the size of the model class, we can effectively “exclude” models that define the same empirical loss across different datasets. Let $\bar{{\boldsymbol{\Theta}}}\subseteq{\boldsymbol{\Theta}}$ be a subset of the model class where, if ${\boldsymbol{\theta}},{\boldsymbol{\theta}}^{\prime}\in\bar{{\boldsymbol{\Theta}}}$, there exists a dataset $D\sim\nu^{n}$ such that $\hat{L}(D,{\boldsymbol{\theta}})\neq\hat{L}(D,{\boldsymbol{\theta}}^{\prime})$. The following result provides a refined bound, based on the size of $\bar{{\boldsymbol{\Theta}}}$.
Corollary 4.2.
$L({\boldsymbol{\theta}})\leq\hat{L}(D,{\boldsymbol{\theta}})+{\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\tfrac{1}{n}\ln\tfrac{|\bar{{\boldsymbol{\Theta}}}|}{\delta}\right)\,.$
For the remainder of this work, we keep the term $k^{p}$ in all of our results. However, in each case, we can replace $k^{p}$ with the typically smaller term $|\bar{{\boldsymbol{\Theta}}}|$ by applying the result above. The following result shows that, when $n$ goes to infinity, the proposed complexity measure converges with rate $1/\sqrt{n}$ to the (scaled) standard deviation of the loss function.
@@
The relevant property of this novel complexity measure is that it is a perfectly tight proxy of the expected loss $L({\boldsymbol{\theta}})$ for models that interpolate the training data, $\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon$, even if the model class is over-parameterized,
@@
Note that in the above result , the event that holds with probability $1-\delta$ is the implication in Equation 3. Using there results, we can derive an algorithm-distribution-dependent bound that is perfectly tight for any data distribution and any learning algorithm; even for over-parameterized model classes. However, for this bound to be tight it is imperative that the model interpolates the training data. Formally, for any fixed $\epsilon>0$, with $2^{\mathcal{X}\times\mathcal{Y}}$ denoting the power set of all possible data sets of any size; let $\mathcal{A}_{\epsilon}:2^{\mathcal{X}\times\mathcal{Y}}\to{\boldsymbol{\Theta}}$ be an algorithm that takes any data set and returns a model ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}$ such that its training loss is lower than $\epsilon$. That is, for any training data set $D\in 2^{\mathcal{X}\times\mathcal{Y}}$, $\mathcal{A}_{\epsilon}$ verifies that ${\hat{L}}({\cal A}_{\epsilon}(D),D)\leq\epsilon$. This kind of algorithms are quite common in machine learning where the model space ${\boldsymbol{\Theta}}$ is big enough that there exists models capable of memorizing random labels (Zhang et al.,, 2017, Theorem 1). Then, for any ${\cal A}_{\epsilon}$ algorithm, with h.p. $1-\delta$ over $D\sim\nu^{n}$,
@@
This provides an answer to the Open Question 4.1 using this newly presented bound:
@@ 4.2 Smoother Interpolators Generalize Better
This result shows the relation between the smoothness and the variance the loss function between a pair of models ${\boldsymbol{\theta}}$ and ${\boldsymbol{\theta}}^{\prime}$. As a result, we can intuitively understand this notion of smoothness as a generalization of the variance of the loss of a model.
Using this definition of smoothness and the results presented in Section 4.1, we can easily derive a result providing an answer to which interpolator generalizes better,
@@
Figure 5: Visualization on the presented notion of smoothness. Two models, a linear model ${\boldsymbol{\theta}}_{1}$ and a complex model ${\boldsymbol{\theta}}_{2}$, are evaluated on two data-generating distributions $\nu_{1}$ (first row) and $\nu_{2}$ (second row). The first column shows the data and models, the second shows rate functions, and the third displays $\hat{L}(D,{\boldsymbol{\theta}})$ distributions.
@@
Figure 5 shows a synthetical example to highlight how the notion of smoothness depends on the specific data-generating distribution, and why one model is smoother than another only relative to such distribution. The example considers two different data generating distributions $\nu_{1}$ (first row, which adds random noise to a linear function) and $\nu_{2}$ (second row, adds random noise to a complex sinusoidal function). The considered loss is the mean squared error. The introduced notion of smoothness using the rate function shows how ${\boldsymbol{\theta}}_{1}$ (a linear model) is smoother than ${\boldsymbol{\theta}}_{2}$ (a more complex model) under $\nu_{1}$. In fact, under $\nu_{1}$, the distribution of $\hat{L}(D,{\boldsymbol{\theta}}_{1})$ is more concentrated around a smaller mean value. However, the second row of this example also shows how the more complex model ${\boldsymbol{\theta}}_{2}$ can be smoother than a linear model ${\boldsymbol{\theta}}_{1}$ under a different data-generating distribution $\nu_{2}$ for exactly the same reasons. With this example we want to highlight that a good notion of smoothness (as the one presented here) should consider the data in which the models are being evaluated, rather than just the complexity of the function they induced.
The results from this section clearly indicate that the generalization error of a interpolator is defined by its level of smoothness and its rate function. The question is whether this theoretical characterizations, which relies on distribution-dependent quantities, is useful for understanding the inner workings of current learning techniques and complex phenomenons appearing in deep learning. In Section 5, we show how the PAC-Chernoff bound and our smoothness criteria are powerful enough to analyze the so-called double-descent phenomenon (Belkin et al.,, 2019). In Section 6, we examine the relationship between the smoothness of a model and widely used regularization techniques such as the parameter norm of a model, distance from initialization, input-gradient norm, and Lipschitz constant. Many of these quantities have also been previously used as measures of smoothness or complexity of a model (Neyshabur et al., 2017a, ). In Section 7, we show why invariant architectures and data-augmentation induce smoother interpolators. Finally, in Section 8, we examine how over-parametrization is a necessary condition for having smoother interpolators. In Section 10, we also elaborate on the need to stablish connections our definition of smoothness with current existing definitions (Keskar et al., 2017b, ; Shalev-Shwartz & Ben-David,, 2014).
5 Understanding Double-Descent with PAC-Chernoff Bounds
The existing literature consistently demonstrates that interpolators with a larger number of parameters tend to perform better. A finding that defies traditional beliefs from classical statistical learning theory. Traditionally, it was assumed that increasing the number of parameters in a model would lead to higher overfitting and, consequently, poor generalization performance. However, this perspective has been challenged by the phenomenon of the double-descent curve (Belkin et al.,, 2019), which exemplifies a shift in dynamics as models enter the interpolation regime. In this regime, an increase in parameters paradoxically leads to improved performance. This counter-intuitive behavior suggests that once models begin to interpolate, their generalization performance improves as they grow larger.
Size Train Acc. Test Acc. Test NLL Bound
$4$k $43.82\%$ $42.74\%$ $1.56$ $2.93$
$32$k $70.84\%$ $59.86\%$ $1.17$ $2.01$
$85$k $86.69\%$ $62.73\%$ $1.35$ $1.77$
$163$k $97.65\%$ $63.97\%$ $1.82$ $1.94$
$266$k $99.84\%$ $65.19\%$ $2.15$ $2.19$
$395$k $100.0\%$ $65.74\%$ $2.28$ $2.30$
$548$k $100.0\%$ $67.27\%$ $2.10$ $2.11$
$727$k $100.0\%$ $69.88\%$ $1.84$ $1.86$
$931$k $100.0\%$ $69.11\%$ $1.83$ $1.84$
$1161$k $100.0\%$ $69.82\%$ $1.72$ $1.74$
Figure 6: Illustration of the double-descent phenomenon: Evolution of the train loss, test loss and the proposed PAC-Chernoff Bound of nested convolutional neural networks with increasing number of parameters. Increasing the number of parameters promotes models with worse population loss until the critical region; further increasing the number of parameters decreases the population error. The proposed bound is perfectly tight for interpolators. The point where models reach the interpolation regime is marked with a black line. The training and test loss of models are marked with points and their smoothed curve is shown. Only a few of the total models are shown in the table.
Figure 6 also displays the evolution of our PAC-Chernoff bound for the whole range of models. As theoretically shown, this bound is perfectly tight for interpolators. In the classical regime, the bound also presents the so-called double descent phenomena, even thought its tightness is not guaranteed. This highlights how distribution-dependent bounds are powerful enough to capture the complex dynamics emerging in the interpolation regime.
Open Question 3 Why does the generalization performance of interpolators improve with an increasing number of parameters?
Figure 7: Rate functions obtained by interpolating CNNs with 4k-1161k parameters on Cifar10. Models correspond to those in the table of Figure 6. Parameters are increased by increasing the number of channels in the architecture, resulting in nested model classes. Two images are shown where models are divided so that the rate function decrease on the left figure and increase on right one.
The complexity term of our own PAC-Chernoff bound ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)$ monotonically increases with the size of the mode class, which, in principle, would contradict the double-descent phenomena on the bound and the fact that generalization error is reduced as we increase the number of models. The following result, viewed as an extension of Theorem 4.6, offers a partial explanation for this puzzling phenomena. More precisely, it demonstrates that if the interpolators become progressively smoother (higher rate function), their generalization error will be reduced, despite being part of a model class characterized by a greater number of parameters.
Theorem 5.1.
Let ${\boldsymbol{\Theta}}\subset{\boldsymbol{\Theta}}^{\prime}$ be two nested model classes with $p<p^{\prime}$ parameters respectively. For any $\epsilon>0$, with h.p. $1-\delta$ over $D\sim\nu^{n}$, for any ${\boldsymbol{\theta}}^{\prime}\in{\boldsymbol{\Theta}}^{\prime}$, ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}$, simultaneously
[…]
$\Downarrow$
Figure 7 (right) illustrates the rate functions of the interpolators from Figure 6, highlighting how larger interpolators become increasingly smoother, meaning their rate functions are progressively higher (i.e., they are progressively smoother). This explains how it is possible to obtain larger interpolators with smaller generalization error. In fact, by reversing the implication of Theorem 5.1, we can infer that if a larger interpolator exhibits better generalization performance, then, it cannot be less smooth than a smaller interpolator.
Corollary 5.2.
Let ${\boldsymbol{\Theta}}\subset{\boldsymbol{\Theta}}^{\prime}$ be two nested model classes with $p<p^{\prime}$ parameters respectively. For any $\epsilon>0$, with h.p. $1-\delta$ over $D\sim\nu^{n}$, for any ${\boldsymbol{\theta}}^{\prime}\in{\boldsymbol{\Theta}}^{\prime}$, ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}$, simultaneously,
$\Downarrow$
Consequently, Open Question 5 can be reformulated in terms of model smoothness rather than generalization performance.
Open Question 4 Why does the smoothness of interpolators improve with an increasing number of parameters?
Based on Theorem 5.1 and Corollary 5.2, addressing Open Question 5 is equivalent to answering Open Question 5. In other words, if we could answer why the rate function of the interpolators is increasingly higher as we increase the number of parameters in nested model classes, Theorem 5.1 will effectively explain why the double descent phenomena, and in particular, why the generalization performance of interpolators enhances as the number of parameters grows. However, to the best of our knowledge, there is no direct answer to this behavior without incurring in additional hypothesis on the matter.
Our rationale for explaining this phenomenon posits that, by increasing the number of parameters, larger neural networks are better at capturing invariances in the data, a notion that has been supported both theoretically (Bronstein et al.,, 2021) and empirically (Goodfellow et al.,, 2009) by the machine learning community. Additionally, as will be shown in Section 7, this capability of capturing invariances leads to smoother models with higher rate functions. Combining these two factors, Theorem 5.1 can help to elucidate why the generalization performance of interpolators improves with an increase in parameters. The experimental setup presented in Figures 6 and 7 support this understanding.
Finally, it is also quite interesting to look at the models on the left part of the interpolation threshold of Figure 6. Surprisingly, the sequence of rate functions, displayed in Figure 7 (left), get increasingly lower as the size of the mode class increases, just the opposite of what happens in Figure 7 (right). In the classical regime, increasing the number of parameters leads to less smooth models with higher generalization error. However, explaining this behavior in the non-interpolation regime is out of the scope of this work, as the presented PAC-Chernoff bound is not tight in that regime and no theoretical guarantees can be derived in such framework. However, it is worth noticing that, even with the lack of tightness guarantees, the proposed bound is close enough to showcase the double descent phenomena itself.
In the following section, we will connect the proposed oracle bound with many existing regularization techniques, allowing to explain, in an unified framework, the effectiveness in terms of generalization.
@@ 6 Explicit Regularization
Structural risk minimization (Shawe-Taylor et al.,, 1998) is a learning principle based on regularizers that penalize the complexity of a model when minimizing the training loss. Let $r({\boldsymbol{\theta}})$ denote a regularizing function, the learning objective is then given by,
@@
High-probability bounds of the form given in Equation (1), $L({\boldsymbol{\theta}})\leq\hat{L}(D,{\boldsymbol{\theta}})+{\cal C}({\boldsymbol{\theta}},D,\delta)$, have been traditionally used to justify the use of regularizers and to derive novel ones. According to these bounds, an obvious choice for regularizes is the complexity term, $r({\boldsymbol{\theta}})={\cal C}({\boldsymbol{\theta}},D,\delta)$. Then, the structural risk minimization problem given in Equation (4) would correspond to minimize a high-probability upper bound over the expected loss $L({\boldsymbol{\theta}})$. Many regularizing methods have been studied from this perspective. One of the most commons one is being the $\ell^{2}$-norm regularization and other variations based on spectral norms, which recurrently show up within the complexity term of many generalization bounds (Bartlett et al.,, 2017). However, as discussed in the introduction, these bounds are known to be loose in the context of over-parameterized model classes interpolating the data. In consequence, the rationale of using their associated complexity measures ${\cal C}({\boldsymbol{\theta}},D,\delta)$ as regularizes is no longer as strong as it used to be in learning setups where the model class is not over-parameterized.
Open Question 5 Is there a regularizer $r({\boldsymbol{\theta}})$ that ensures near-optimal performance for over-parameterized model classes interpolating the training data?
The following result provides an answer to Open Question 6. It shows that the complexity measure of the PAC-Chernoff bound given in Theorem 4.1 can be used as a regularizer. That is to say, using the inverse rate function ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)$, we get an optimal regularizer for over-parameterized interpolators. More precisely, for any $\epsilon>0$, let ${\boldsymbol{\theta}}^{\star}_{\epsilon}$ and ${\boldsymbol{\theta}}^{\crossproduct}_{\epsilon}$ denote the interpolator with the best generalization performance and the interpolator with the smallest inverse rate function, respectively,
@@
In simpler terms, favoring models with a small inverse rate function is the same as favoring models with a larger rate function, which essentially means choosing smoother models. This finding highlights that the smoothest interpolator not only fits the data but also performs nearly optimally when it comes to generalization.